### v1

@@ 6.2 Effect of Confidence Thresholds on MBTI Classification
Adopting higher confidence thresholds leads to a reduction in the number of classified data points, highlighting a trade-off between accuracy and data coverage. At the highest threshold of 0.99, approximately 25% of the total dataset remains classified. This suggests that although RoBERTa’s predictions exhibit high accuracy at increased confidence levels, a significant amount of data remains unclassified.
Notwithstanding this limitation, our approach provides significant practical value. Adjusting the confidence threshold provides practitioners with flexibility, enabling them to optimize the balance between broad coverage and precision according to their specific application requirements. Settings that require high certainty, such as psychological assessment, recruitment, or personalized recommendation systems, would significantly benefit from prioritizing precision over comprehensive coverage.
Our results demonstrate the substantial benefits of confidence-threshold tuning in transformer-based implicit personality classification tasks, emphasizing the methodological rigor and practical relevance of our approach.
A significant avenue for future research involves utilizing multiple messages per user instead of focusing on single messages, which may enhance both prediction accuracy and coverage through aggregation. Considering multiple conversational samples per individual allows models to address the coverage limitation seen at high confidence thresholds, thereby improving the practical applicability and robustness of implicit personality detection from conversational data.

### v2

@@ 6.2 Effect of Confidence Thresholds on MBTI Classification
Table 2 presents the relationship between different confidence thresholds, classification accuracy, and the proportion of classified data. As shown, increasing the confidence threshold leads to a notable increase in classification accuracy. However, this comes at the cost of reduced data coverage. For example, at a threshold of 0.50, the model achieves 52.83% accuracy while covering 85% of the dataset, whereas at the highest threshold of 0.99, accuracy reaches 86.16% but covers only 26% of the dataset.
Table 2: Effect of Confidence Thresholds on MBTI Classification
Threshold Accuracy (%) Data Coverage (%)
0.50 52.83 85
0.60 55.51 77
0.70 59.82 68
0.80 64.12 59
0.90 70.16 48
0.99 86.16 26
@@
Adopting higher confidence thresholds leads to a reduction in the number of classified data points, highlighting a trade-off between accuracy and data coverage. As seen in Table 2, at the highest threshold of 0.99, approximately 26% of the total dataset remains classified. This suggests that although RoBERTa’s predictions exhibit high accuracy at increased confidence levels, a significant amount of data remains unclassified.
Nevertheless, this limitation is offset by the immense value that such a highly accurate classification provides. The ability to predict an individual’s MBTI personality type with 86.16% accuracy, solely based on their text messages, is a remarkable achievement in the field of computational psychology and AI-driven personality analysis. This level of precision suggests that even in real-world applications where only a fraction of messages can be confidently classified, the insights generated by the model remain highly reliable and valuable.
Such a high-confidence personality prediction system can be particularly useful in applications where accurate personality profiling is essential, such as in psychological assessments, AI-driven recruitment systems, and adaptive learning platforms that personalize educational content based on inferred personality traits. The ability to infer personality traits implicitly, without requiring users to take explicit psychological tests, not only enhances user experience but also opens the door to more seamless and non-intrusive applications in personalized AI systems.
Furthermore, our results demonstrate the substantial benefits of confidence-threshold tuning in transformer-based implicit personality classification tasks, emphasizing the methodological rigor and practical relevance of our approach. By dynamically adjusting the confidence threshold, practitioners can optimize their models for either broad coverage or high precision, depending on their specific application requirements.
A significant avenue for future research involves utilizing multiple messages per user instead of focusing on single messages, which may enhance both prediction accuracy and coverage through aggregation. Considering multiple conversational samples per individual allows models to address the coverage limitation seen at high confidence thresholds, thereby improving the practical applicability and robustness of implicit personality detection from conversational data. By leveraging larger conversational contexts, it is likely that we can push classification accuracy even further while maintaining broader coverage.